OpenStudy (anonymous):

suppose that a,b are elements of a group and a^5=e, b^4=e, ab=ba^3 Prove that (i) a^b =ba (ii) ab^3=b^3a^2

OpenStudy (perl):

you sure there are no typos, for part a) a^b = ba ?

OpenStudy (perl):

so this is like a puzzle

OpenStudy (perl):

apply multiplication on both sides

OpenStudy (anonymous):

dont understand

OpenStudy (perl):

ab=ba^3 multiply both sides by a^2 (on the right) ab *a^2 = b*a^3 * a^2 aba^2 = b* a^5 aba^2 = b*e aba^2 = b

OpenStudy (perl):

ab=ba^3 multiply both sides by a^2 (on the right) ab *a^2 = b*a^3 * a^2 ab*a^2 = b* a^5 aba^2 = b*e aba^2 = b

OpenStudy (perl):

i think you can make a table

OpenStudy (perl):

i will be able to answer more quickly if i see the original worksheet or problem

OpenStudy (perl):

can you take a picture of the question and upload it?

OpenStudy (anonymous):

i cant it was written down on a piece of paper